3 5/6 - 1 1/4 + 1 1/2
Found 2 solutions by Alan3354, mathsolverplus:
Answer by Alan3354(69443)

3 5/6 - 1 1/4 + 1 1/2
===========
Change to "improper fractions."
= (23/6) - (5/4) + (3/2)
You might spot the the LCD (Least Commom Denominator) is 12.
You might not.
-----------
The product of the DENs always works.
6*4*2 = 48
---
(23/6) - (5/4) + (3/2) = (184/48) - 60/48) + (72/48)
= (184 - 60 + 72)/48
= 196/48
Then reducing is easy.
= 49/12
========================
You can use the product of the DEN's, then, if it's asked for, the result is in the LCD.
-----
Wasting time and fretting about the LCD accomplishes nothing.
Using the product ALWAYS works, and reducing is easy.

Answer by mathsolverplus(88)

The least common multiple factor among 6,4, and 2 is 12.
3 (10/12) - 1 (3/12) + 1 (6/12)
= 3-1+1 ((10-3+6)/12))
= 3 (13/12)
=4 (1/12)



If you need additional help, please follow us on Instagram or message us via social media! We also help with SAT Math Questions!

***Instagram (@mathsolver.plus): https://www.instagram.com/mathsolver.plus/
Twitter (@MathSolverPlus): https://twitter.com/MathSolverPlus
Snapchat (@mathsolverplus)
Email: mathsolverplus@gmail.com